OpenStudy (anonymous):

Check for extraneous solutions. 1.\[\sqrt{x-3}=x-3\] 2.\[\sqrt{5x-9}=x-9\]

OpenStudy (anonymous):

please show the work

OpenStudy (anonymous):

square both sides. you will get a quadratic equation and then solve the quadratic

OpenStudy (anonymous):

\[\sqrt{x-3}=x-3 \rightarrow x-3 = (x-3)^{2}\rightarrow x-3 = x ^{2}-6x+9\]

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

\[x ^{2}-7x+12 =0 \rightarrow x=3,4\]

OpenStudy (anonymous):

x^2-7x+12=0 \[(x-3)(x-4)=0\]

OpenStudy (anonymous):

the second problem is the same as the first problem, with different numbers. follow the same procedure

OpenStudy (anonymous):

5x-9=x^2-18x+81 x^2-23x+90=0 (x-18)(x-5)=0 x=18,5